OpenStudy (mowrey):

If you have two more than 3x the number of cards that someone has. And if you give that someone 12 of your cards, then you each will have the same number of cards. How many cards do you have?

OpenStudy (mowrey):

Here's my take on this problem: Guess: 12 Check: (12 * 3) + 2 = 36 + 2 = 38; 38 - 12 = 26; 26 - 12 = 14 Too High Guess: 10 Check: (10 * 3) + 2 = 32; 32 - 12 = 20; 20 - 12 = 8 Too Low Guess: 11 Check: (11 * 3) + 2 = 35; 35 - 12 = 23 Correct, since 23 - 11 = 12 (two cards more than the number of cards Antonio has)